# Non-linear shattered pellet injection modelling in ASDEX Upgrade

### v1

@@ 3.4 Effect of penetration speed
To further study the effect of penetration speed on disruption mitigation, simulations are carried out based on the cases LF_HV_Ne10 and MF_HV_Ne10 as described in Tab. 1. Generally speaking, the shard configuration is determined by two main factors: the impact speed ($v_{\perp}$), which is the pellet velocity perpendicular to the shattering plane and influences the size of the shards, and the penetration speed ($v_{\parallel}$), which is the pellet velocity parallel to the shattering plane and affects the velocity distribution of the shards after shattering. To isolate the effects of these two factors, we maintain the same shard numbers and sizes as in the cases LF_FV_Ne10 and MF_FV_Ne10 while reducing the velocities of the shards to half of their original values (Half Velocity). This approach retains the impact speed while altering the penetration speed.
@@
Figure 15: Temporal evolution of (a) total thermal energy, (b) plasma current, (c) total assimilated neon quantity and (d) radiation power for the 53 and 199 shard cases with different penetration speed and 10% neon. The traces of the two slower cases are plotted with the time shifted to align the arrival time, indicated by $t_{arr}$, of the shards with that of the faster cases.
@@
Figure 16: Flux averaged temperature profile evolution for different penetration speed with 10% neon. The dashed lines indicate the estimated shard trajectory.
@@
Figure 17: (a) Plasma current $I_{p}$ and (b) thermal energy $W_{th}$ evolution for the case LF_HV_Ne10 in comparison with the experiment discharge #40673. In this particular experimental discharge, the pellet speed is 221 m/s, the shatter angle is 25 degrees, and the neon fraction is 10%. The traces of the simulation are plotted with the time shifted to align the arrival time of the shards with experiment, which is annotated by the cyan dotted line. The orange and blue dotted lines indicate the time of $I_{p}$ dip for the two cases, respectively.
@@
The macroscopic quantities for different penetration speeds are compared in figure 15. The traces of the two slower cases are plotted with the time shifted to align the arrival time of the shards with that of the faster cases, for a clearer comparison. Numerical results suggest that faster penetration speeds can facilitate the material assimilation, especially during the early stage after injection, leading to shorter TQ duration. With a faster penetration speed, the shards reach the hotter regions more quickly enhancing stronger ablation, and hence increasing the impurity radiation. The initial thermal energy drop is also more steep, comparing the slow speed cases. The transition from an inside-out to an outside-in cooling pattern can be observed with decreasing penetration speed, as shown in figure 16. Due to the slower penetration speed, the shards remain longer at the edge during their penetration, allowing for more efficient impurity assimilation at the edge, which contributes to the formation of the cold front.
@@
As the cold front forms at the edge, a more pronounced $I_{p}$ spike is observed in the simulation. Figure 17 presents the plasma current and thermal energy evolution for the case LF_HV_Ne10, compared with the experiment discharge #40673. A comparable $I_{p}$ spike with experiment is observed in the simulation. On the other hand, the $I_{p}$ dip in the simulation occurs significantly earlier than that in experiment. This earlier dip is attributed to the excessively sharp thermal energy drop following the shard arrival, as shown in the thermal energy evolution. This rapid energy loss leads to an unrealistically short pre-TQ duration, as discussed in section 3.2.
@@ 3.5 General figures of merit
The aforementioned simulations provide valuable insights into the plasma response to various shard parameters. In this subsection, general trends and summary plots are presented, highlighting the dependence on these parameters and how the results align with experimental observations.
@@
Figure 20: The total radiation power vs. toroidal peaking factor. The TPF is calculated by $P_{rad}^{max}/\Sigma(P_{rad,i}+P_{rad,i+1})/2*\Delta_{\phi}(i,i+1)/360$, where $P_{rad}^{max}$ is the maximum radiation among the 5 sectors installed with foil bolometer, $(P_{rad,i}+P_{rad,i+1})/2$ is the average radiation between the neighboring sectors and $\Delta_{\phi}(i,i+1)$ is the toroidal angle between the two neighboring sectors. The contour plot represents the constant values of the product TPF * $P_{rad}$.
@@
Figure 22: Pre-TQ+TQ duration, period from shard arrival to the end of TQ, for all neon doped cases. For the pure deuterium case, since a complete TQ does not occur, this duration is not available.
@@
Figure 18 illustrates the radiation fraction similar to the experimental definition following Ref. [20] $f_{rad}=W_{rad}/(W_{th}(t=0)+W_{ohm})$ at the time point when the plasma current $I_{p}$ has dropped to 80% of its original value for a consistent comparison, since not all the cases reach at the same CQ level by the end of simulation. Here, $W_{rad}$ is the total radiated energy, $W_{th}(t=0)$ is the thermal energy at the beginning, and $W_{ohm}$ is the total ohmic heating energy in the considered time interval. It should be mentioned that, in our current modeling, as described in section 2, the deuterium neutral fluid and neutral line radiation are not considered. Additionally, no background impurities are included in the 3D simulations. Consequently, for the pure deuterium case, the radiation fraction is 0%. As for neon-doped pellets, the radiated energy accounts for approximately 40% of the lost energy for the 0.12% neon case, and 45% for the 1% neon case. Furthermore, the size of the shards does not have any significant impact on the radiation fraction in the 0.12% neon cases, owing to the low radiative cooling effect during the first stage of thermal energy loss. For 10% neon cases, the radiation fraction ranges from 56% - 74%, depending on different injection scenarios. Faster and smaller shards exhibit better performance in terms of radiation fraction, as they generate higher radiation during the first stage of thermal energy loss.
@@
To compare with experiments, the radiation fraction, calculated using the above definition at the end of each simulation, is plotted alongside experimental data in figure 19. The experimental data is derived until the end of the CQ [20]. The overall radiation fraction in the simulations is relatively lower than that observed in experiments. This discrepancy is primarily due to excessive thermal energy loss during the first stage of the cooling process, which is characterized by a low radiation level. The key to improve the radiation fraction lies in increasing radiative cooling during the first stage. Notably, the case with 10% neon and 1105 shards shows a comparable radiation fraction to experimental results, owing to more effective radiative coverage in the initial cooling phase.
@@
The total radiation power versus the toroidal peaking factor (TPF) is plotted in figure 20, where the TPF is obtained using the same formula in experiments [40]. The key message here is that there are two distinct radiation peaks, which are much more prominent in the large neon fraction cases. The first peak occurs at the initial injection and the second at the end of the TQ. The first peak, characterized by a relatively large TPF, is mainly caused by the outward heat flux that result in sharp ablation of the shards during the first stage of thermal energy drop, which is overestimated in simulations and not observed in experiments. The smaller shard case exhibits a stronger TPF and higher radiation at this stage. The TPF at the TQ is generally very small for all cases, as the ablated material has spread spatially at the latter phase, which is also observed experimentally. It is crucial to prevent high toroidal peaking during the initial injection phase in experiments, as simulations suggest that this phase is more likely to produce a high TPF, despite the fact that our current simulations tend to overestimate the TPF at this stage. One possible explanation for the more peaked early-phase radiation peaking observed in our simulations is opacity. During the early phase, the local density around the shards could become very high so that the plasma is not transparent for line radiation, reducing the local radiation power density around the shards in reality [41, 42].
@@
In addition, the distribution of radiation across different toroidal sectors in the experiment is more uniform (except S9) compared to what observed in simulation, as illustrated in figure 21. In the simulation, the difference in radiation between sectors near the injection plane and those far from it is more significant, comparing with experiments [40]. This may be due to the fact that the equilibrium toroidal flow is not considered in the simulation.
@@
Figure 22 presents the pre-TQ+TQ duration for all neon-doped cases. In the case of pure deuterium injection, no complete TQ is observed due to the lack of radiative cooling, resulting in an effectively infinite TQ duration. This behavior aligns with non-disruptive shots or those with extended pre-TQ durations seen in experiments following pure deuterium injection [20]. For neon-doped injections, even a slight neon fraction of 0.12% reduces the pre-TQ+TQ time to 3 – 5 ms. As the neon fraction increases to 10%, this duration further decreases to approximately 1 ms. Additionally, for cases with low neon fractions, the pre-TQ+TQ time shows a noticeable dependence on shard size, which can be attributed to the significant differences in material assimilation in our simulations.
@@
In terms of the assimilation, figure 24 exhibits the assimilated neon atoms for all the neon doped cases. As expected, neon assimilation in absolute numbers generally increases with higher neon fractions. For the 0.12% neon cases, there is a clear dependence on the shard size, with smaller shards showing higher assimilation, which deviates from the experimental results. This is because our current simulation fails in accurately simulating the plasmoid drift and does not take into account the rocket effect, as we discussed in section 3.3. On the other hand, for cases with higher neon fractions, such as the 10% neon case, the dependence on shard size is less pronounced due to the extended ablation duration for larger shard case. This part of results is consistent with experimental observation, presumably due to the suppression of plasmoid drift at high neon fraction in experiments. With respect to penetration speed, faster shards result in higher assimilation during the early stage after injection, as shown in figure 15; however, the difference is very small in the end.
Regarding the total assimilation, figure 25 displays the total assimilated atoms (neon + deuterium) for all cases. The trend reverses when considering the total assimilated atoms, with total assimilation increasing as the neon fraction decreases. This is due to the lower assimilation ratio observed in high neon fraction cases, primarily because of the much shorter TQ duration. As these pellets contain comparable amounts of deuterium, and deuterium atoms far outnumber neon atoms within the pellet, a higher ablation ratio leads to enhanced deuterium assimilation. The dependence on shard size and speed follows a similar pattern to that of neon assimilation.
@@ 4 Summary and discussion
In this study, 3D non-linear SPI simulations for an AUG H-mode scenario are carried out using realistic parameters with the 3D non-linear JOREK code, aimed at validating optimal shard configurations for the ITER DMS. The key findings can be summarized as follows:
@@
The total assimilation becomes less efficient with increasing neon content. Higher neon fraction can lead to shorter ablation duration, resulting in a lower ablation ratio. For the low neon fraction (0.12%) cases, the assimilation exhibits a strong dependence on shard size. Smaller shards increase the surface area available for ablation, leading to higher assimilation. This trend diverges from experimental observations, where larger shards typically lead to higher material assimilation. This discrepancy arises from inaccuracies in simulating the pre-TQ phase. A larger number of smaller shards can induce stronger plasmoid drift due to quicker ablation and larger pressure peaking, which is not accurately captured in the current simulation. Additionally, the current model does not incorporate the rocket effect, which might be particularly important for small shards, as their smaller mass makes them more susceptible to this effect. For the cases with high neon fraction (10%), the dependence of assimilation on shard size is less pronounced. The smaller shards lead to shorter ablation duration, even though the ablation rate is higher. This part of results aligns with experimental observations, likely due to the suppression of plasmoid drift at high neon fraction injections in experiments.
@@
In cases involving large shards (53 shards), an inside-out cooling pattern emerges due to less efficient edge cooling. The results suggest that smaller and slower shards promote edge cooling, which aids in forming a cold front and shifting the cooling pattern from inside-out to outside-in. The plume generated by smaller shards is more elongated, ensuring a continuous presence of shards at the edge. Additionally, smaller shards demonstrate more efficient assimilation during the initial stages, while slower shards remain longer at the edge during penetration, further enhancing edge cooling. As the cold front forms at the edge, an $I_{p}$ spike comparable to the experimental one is observed in the simulation.
@@
The overall radiation fractions observed in the simulations are slightly lower than those seen in experiments, likely due to overestimated thermal energy loss in the first stage of cooling. This discrepancy is primarily due to excessive thermal energy loss during the first stage of cooling, which is not sufficiently radiated. As for the radiation asymmetry, two radiation peaks are observed in large neon fraction cases, one at initial injection and another at the end of the TQ. The first peak, caused by outward heat flux during the initial thermal energy drop, is overestimated in our simulations. The smaller shard case shows stronger peaking factors and higher radiation during this stage. The radiation peak at TQ shows a small peaking factor due to a more relaxed plasma state. The radiation distribution in experiments is more uniform across toroidal sectors than in simulations, likely because our current simulations do not include the equilibrium toroidal flow and the opacity effect. The current work aims to validate the SPI model in JOREK and compare the general trends with experimental results. More detailed comparisons with experiments will be carried out in the future work.

### v2

@@ 3.4 Effect of penetration speed
To further study the effect of penetration speed on disruption mitigation, simulations are carried out based on the cases LF_HV_Ne10 and MF_HV_Ne10 as described in Tab. 1. Generally speaking, the fragment configuration is determined by two main factors: the impact speed ($v_{\perp}$), which is the pellet velocity perpendicular to the shattering plane and influences the size of the fragments, and the penetration speed ($v_{\parallel}$), which is the pellet velocity parallel to the shattering plane and affects the velocity distribution of the fragments after shattering. To isolate the effects of these two factors, we maintain the same fragment numbers and sizes as in the cases LF_FV_Ne10 and MF_FV_Ne10 while reducing the velocities of the fragments to half of their original values (Half Velocity). This approach retains the impact speed while altering the penetration speed.
@@
Figure 18: Temporal evolution of (a) total thermal energy, (b) plasma current, (c) total assimilated neon quantity and (d) radiation power for the 53 and 199 fragment cases with different penetration speed and 10% neon. The traces of the two slower cases are plotted with the time shifted to align the arrival time, indicated by $t_{arr}$, of the fragment with that of the faster cases.
@@
Figure 19: Flux-averaged temperature profile evolution for different penetration speed with 10% neon. The dashed lines indicate the estimated trajectory of the bulk fragments.
@@
Figure 20: (a) Plasma current $I_{p}$ and (b) thermal energy $W_{th}$ evolution for the case LF_HV_Ne10 in comparison with the experiment discharge #40673. In this particular experimental discharge, the pellet speed is 221 m/s, the shatter angle is 25 degrees, and the neon fraction is 10%. The traces of the simulation are plotted with the time shifted to align the arrival time of the fragments with experiment, which is annotated by the cyan dotted line. The orange and blue dotted lines indicate the time of $I_{p}$ dip for the two cases, respectively.
@@
The macroscopic quantities for different penetration speeds are compared in figure 18. The traces of the two slower cases are plotted with the time shifted to align the arrival time of the fragments with that of the faster cases, for a clearer comparison. Numerical results suggest that faster penetration speeds can facilitate the material assimilation, especially during the early stage after injection, leading to shorter TQ duration. With a faster penetration speed, the fragments reach the hotter regions more quickly enhancing stronger ablation, and hence increasing the impurity radiation. The initial thermal energy drop is also more steep, comparing the slow speed cases. The transition from an inside-out to an outside-in cooling pattern can be observed with decreasing penetration speed, as shown in figure 19. Due to the slower penetration speed, the fragments remain longer at the edge during their penetration, allowing for more efficient impurity assimilation at the edge, which contributes to the formation of the cold front.
@@
As the cold front forms at the edge, a more pronounced $I_{p}$ spike is observed in the simulation. Figure 20 presents the plasma current and thermal energy evolution for the case LF_HV_Ne10, compared with the experiment discharge #40673. A comparable $I_{p}$ spike with experiment is observed in the simulation. On the other hand, the $I_{p}$ dip in the simulation occurs significantly earlier than that in experiment. This earlier dip is attributed to the excessively sharp thermal energy drop following the fragment arrival, as shown in the thermal energy evolution. This rapid energy loss leads to an unrealistically short pre-TQ duration, as discussed in section 3.2.
@@ 3.5 General figures of merit
The aforementioned simulations provide valuable insights into the plasma response to various fragment parameters. In this subsection, general trends and summary plots are presented, highlighting the dependence on these parameters and how the results align with experimental observations.
@@
Figure 23: The total radiation power vs. toroidal peaking factor for (a) different neon fraction, (b) 10% neon with different fragment size, (c) 0.12% neon with different fragment size and (d) different penetration speed. The TPF is calculated by $P_{rad}^{max}/\Sigma(P_{rad,i}+P_{rad,i+1})/2*\Delta_{\phi}(i,i+1)/360$, where $P_{rad}^{max}$ is the maximum radiation among the 5 sectors installed with foil bolometer, $(P_{rad,i}+P_{rad,i+1})/2$ is the average radiation between the neighboring sectors and $\Delta_{\phi}(i,i+1)$ is the toroidal angle between the two neighboring sectors. The contour plot represents the constant values of the product TPF * $P_{rad}$.
@@
Figure 25: Pre-TQ+TQ duration, period from fragment arrival to the end of TQ, for all neon doped cases. For the pure deuterium case, since a complete TQ does not occur, this duration is not available.
@@
Figure 21 illustrates the radiation fraction similar to the experimental definition following Ref. [20] $f_{rad}=W_{rad}/(W_{th}(t=0)+W_{ohm})$ at the time point when the plasma current $I_{p}$ has dropped to 80% of its original value for a consistent comparison, since not all the cases reach at the same CQ level by the end of simulation. Here, $W_{rad}$ is the total radiated energy, $W_{th}(t=0)$ is the thermal energy at the beginning, and $W_{ohm}$ is the total ohmic heating energy in the considered time interval. It should be mentioned that, in our current modeling, as described in section 2, the deuterium neutral fluid and neutral line radiation are not considered. Additionally, no background impurities are included in the 3D simulations. Consequently, for the pure deuterium case, the radiation fraction is 0%. As for neon-doped pellets, the radiated energy accounts for approximately 40% of the lost energy for the 0.12% neon case, and 45% for the 1% neon case. Furthermore, the size of the fragments does not have any significant impact on the radiation fraction in the 0.12% neon cases, owing to the low radiative cooling effect during the first stage of thermal energy loss. For 10% neon cases, the radiation fraction ranges from 56% - 74%, depending on different injection scenarios. Faster and smaller fragments exhibit better performance in terms of radiation fraction, as they generate higher radiation during the first stage of thermal energy loss.
@@
To compare with experiments, the radiation fraction, calculated using the above definition at the end of each simulation, is plotted alongside experimental data in figure 22. The experimental data is derived until the end of the CQ [20]. The overall radiation fraction in the simulations is lower than that observed in experiments. This discrepancy is primarily due to excessive thermal energy loss during the first stage of the cooling process, which is characterized by a low radiation level. The key to improve the radiation fraction lies in increasing radiative cooling during the first stage. Notably, the case with 10% neon and 1105 fragments shows a comparable radiation fraction to experimental results, owing to more effective radiative coverage in the initial cooling phase.
@@
The total radiation power versus the toroidal peaking factor (TPF) is plotted in figure 23, where the TPF is obtained using the same formula in experiments [42]. The key message here is that there are two distinct radiation peaks, which are much more prominent in the large neon fraction cases. The first peak occurs at the initial injection and the second at the end of the TQ. The first peak, characterized by a relatively large TPF, is mainly caused by the outward heat flux that result in sharp ablation of the fragments during the first stage of thermal energy drop, which is overestimated in simulations and not observed in experiments. The smaller fragment case exhibits a stronger TPF and higher radiation at this stage. The TPF at the TQ is generally very small for all cases, as the ablated material has spread spatially at the latter phase, which is also observed experimentally. It is crucial to prevent high toroidal peaking during the initial injection phase in experiments, as simulations suggest that this phase is more likely to produce a high TPF, despite the fact that our current simulations tend to overestimate the TPF at this stage. One possible explanation for the more peaked early-phase radiation peaking observed in our simulations is opacity. During the early phase, the local density around the fragments could become very high so that the plasma is not transparent for line radiation, reducing the local radiation power density around the fragments in reality [43, 44].
@@
In addition, the distribution of radiation across different toroidal sectors in the experiment is more uniform (except S9) compared to what was observed in simulation, as illustrated in figure 24. In the simulation, the difference in radiation between sectors near the injection plane and those far from it is more significant, comparing with experiments [42]. This may be due to the fact that the equilibrium toroidal flow is not considered in the simulation.
@@
Figure 25 presents the pre-TQ+TQ duration for all neon-doped cases. In the case of pure deuterium injection, no complete TQ is observed due to the lack of radiative cooling, resulting in an effectively infinite TQ duration. This behavior aligns with non-disruptive shots or those with extended pre-TQ durations seen in experiments following pure deuterium injection [20]. For neon-doped injections, even a slight neon fraction of 0.12% reduces the pre-TQ+TQ time to 3 – 5 ms. As the neon fraction increases to 10%, this duration further decreases to approximately 1 ms. Additionally, for cases with low neon fractions, the pre-TQ+TQ time shows a noticeable dependence on fragment size, which can be attributed to the significant differences in material assimilation in our simulations.
@@
In terms of the assimilation, figure 27 exhibits the assimilated neon atoms for all the neon doped cases. As expected, neon assimilation in absolute numbers generally increases with higher neon fractions. For the 0.12% neon cases, there is a clear dependence on the fragment size, with smaller fragments showing higher assimilation, which deviates from the experimental results. This is because our current simulation fails in accurately simulating the plasmoid drift and does not take into account the rocket effect, as we discussed in section 3.3. On the other hand, for cases with higher neon fractions, such as the 10% neon case, the dependence on fragment size is less pronounced due to the extended ablation duration for larger fragment case. This part of results is consistent with experimental observation, presumably due to the suppression of plasmoid drift at high neon fraction in experiments. With respect to penetration speed, faster fragments result in higher assimilation during the early stage after injection, as shown in figure 18; however, the difference is very small in the end.
Regarding the total assimilation, figure 28 displays the total assimilated atoms (neon + deuterium) for all cases. The trend reverses when considering the total assimilated atoms, with total assimilation increasing as the neon fraction decreases. This is due to the lower assimilation ratio observed in high neon fraction cases, primarily because of the much shorter TQ duration. As these pellets contain comparable amounts of deuterium, and deuterium atoms far outnumber neon atoms within the pellet, a higher ablation ratio leads to enhanced deuterium assimilation. The dependence on fragment size and speed follows a similar pattern to that of neon assimilation.
@@ 4 Summary and discussion
In this study, 3D non-linear SPI simulations for an AUG H-mode scenario are carried out using realistic parameters with the 3D non-linear JOREK code, aimed at validating optimal fragment configurations for the ITER DMS. The key findings can be summarized as follows:
@@
The total assimilation becomes less efficient with increasing neon content. Higher neon fraction can lead to shorter ablation duration, resulting in a lower ablation ratio. For the low neon fraction (0.12%) cases, the assimilation exhibits a strong dependence on fragment size. Smaller fragments increase the surface area available for ablation, leading to higher assimilation. This trend diverges from experimental observations, where larger fragments typically lead to higher material assimilation. This discrepancy may stem from inaccuracies in simulating the pre-TQ phase. It is possible that a larger number of smaller fragments could enhance plasmoid drift due to quicker ablation and increased local pressure peaking. Additionally, the current model does not incorporate the rocket effect, which might be particularly important for small fragments, as their smaller mass makes them more susceptible to this effect. For the cases with high neon fraction (10%), the dependence of assimilation on fragment size is less pronounced. The smaller fragments lead to shorter ablation duration, even though the ablation rate is higher. This part of results aligns with experimental observations, likely due to the suppression of plasmoid drift at high neon fraction injections in experiments.
@@
In cases involving large fragments (53 fragments), an inside-out cooling pattern emerges due to less efficient edge cooling. The results suggest that smaller and slower fragments promote edge cooling, which aids in forming a cold front and shifting the cooling pattern from inside-out to outside-in. The plume generated by smaller fragments is more elongated, ensuring a continuous presence of fragments at the edge. Additionally, smaller fragments demonstrate more efficient assimilation during the initial stages, while slower fragments remain longer at the edge during penetration, further enhancing edge cooling. As the cold front forms at the edge, an $I_{p}$ spike comparable to the experimental one is observed in the simulation.
@@
The overall radiation fractions observed in the simulations are slightly lower than those seen in experiments, likely due to overestimated thermal energy loss in the first stage of cooling. This discrepancy is primarily due to excessive thermal energy loss during the first stage of cooling, which is not sufficiently radiated. This is likely due to the overestimation of the parallel thermal conduction in our simulations. There are two major ways in which an overestimated thermal conduction can artificially enhance MHD activity in the simulation. First, excessive thermal conduction promotes stronger cooling along magnetic field lines, which leads to a rapid increase in plasma resistivity. This enhanced resistivity, in turn, facilitates the growth of resistive MHD instabilities, thereby exaggerating the MHD response. Secondly, in regions where the magnetic topology becomes stochastic, the overestimated thermal conduction results in unrealistically high thermal energy losses, especially in the initial cooling phase. This excessive energy depletion leads to a shorter and more abrupt pre-TQ phase than what is observed in experiments. As for the radiation asymmetry, two radiation peaks are observed in large neon fraction cases, one at initial injection and another at the end of the TQ. The first peak is caused by the outward heat flux during the initial thermal energy drop. The smaller fragment case shows stronger peaking factors and higher radiation during this stage. The radiation peak at TQ shows a small peaking factor due to a more relaxed plasma state. The radiation distribution in experiments is more uniform across toroidal sectors than in simulations, likely because our current simulations do not include the equilibrium toroidal flow and the opacity effect. The current work aims to validate the SPI model in JOREK and compare the general trends with experimental results. More detailed comparisons with experiments will be carried out in the future work.